OpenStudy (anonymous):

limit-->4 [[(x)^(1/2)]-2]/(x-4) =?(working and hints required)

jhonyy9 (jhonyy9):

when x=4 you see that the denominator is zero - you can write (radical x) -2 ----------- = x-4

OpenStudy (anonymous):

but that isnt the answer,we should not get it in the 0/0 form.it has to be solved algebrecally,but im not sure how.

jhonyy9 (jhonyy9):

ok but hence when x ---> to 4 you can seeing that the function ---> 0 because radical 4 =2 and hence in nominator you will get 2-2 what is equal zero too like the denominator what is zero when x=4

OpenStudy (anonymous):

\[\lim_{x\rightarrow 4} \frac{\sqrt{x}-2}{x-4}=\lim_{x\rightarrow 4}\frac{\sqrt{x}-2}{(\sqrt{x}-2)(\sqrt{x}+2)}=\ldots\]